Calculate the instantaneous power of an ac circuit that has a maximum voltage of 240v and a maximum current of 0.24A with a phase angle of 30 degrees.


1
Expert's answer
2021-06-17T15:29:08-0400


Gives

"I_m=0.24A"

"V_m=240V"

"V_{RMS}=\frac{V_m}{\sqrt2}"

"I_{RMS}=\frac{I_{m}}{\sqrt2}"

"P_{ac}=I_{RMS}V_{RMS}cos\phi"

"P_{ac}=\frac{I_m V_m}{2}cos\phi"

"P_{ac}=\frac{0.24\times240}{2}cos30°"

"P_{ac}=24.94W"